# On the Consistency of Multilingual Context Utilization in Retrieval-Augmented Generation

### v2

@@ Abstract
Retrieval-augmented generation (RAG) with large language models (LLMs) has demonstrated strong performance in multilingual question-answering (QA) tasks by leveraging relevant passages retrieved from corpora. In multilingual RAG (mRAG), the retrieved passages can be written in languages other than that of the query entered by the user, making it challenging for LLMs to effectively utilize the provided information. Recent research suggests that retrieving passages from multilingual corpora can improve RAG performance, particularly for low-resource languages. However, the extent to which LLMs can leverage different kinds of multilingual contexts to generate accurate answers, independently from retrieval quality, remains understudied. In this paper, we conduct an extensive assessment of LLMs’ ability to (i) make consistent use of a relevant passage regardless of its language, (ii) respond in the expected language, and (iii) focus on the relevant passage even when multiple ‘distracting’ passages in different languages are provided in the context. Our experiments with four LLMs across three QA datasets covering a total of 48 languages reveal a surprising ability of LLMs to extract the relevant information from out-language passages, but a much weaker ability to formulate a full answer in the correct language. Our analysis, based on both accuracy and feature attribution techniques, further shows that distracting passages negatively impact answer quality regardless of their language. However, distractors in the query language exert a slightly stronger influence. Taken together, our findings deepen the understanding of how LLMs utilize context in mRAG systems, providing directions for future improvements. All codes and data released at https://github.com/Betswish/mRAG-Context-Consistency.
@@ 1 Introduction
Retrieval-augmented generation has shown strong results in multilingual question-answering (QA) tasks (Chirkova et al., 2024; Thakur et al., 2024). Given a query in the user language, informative passages are retrieved from a reference corpus and provided jointly with the query, promoting the large language model (LLM) to generate more precise responses (Lewis et al., 2020; Asai et al., 2021). In multilingual RAG (mRAG), retrieval can be performed either monolingually or cross-lingually (cf. Figure 1, right). In the former, retrieval is performed only over passages in the same language as the query (Asai et al., 2023; Gao et al., 2023; Fan et al., 2024), which can be successful for high-resource languages. However, this approach is marginally useful, or even harmful, when the question is posed in a low-resource language, since relevant information is likely to be available only in different languages (Muller et al., 2023). In addition, for questions regarding a specific geographical region or culture, essential information may be present only in corpora of the languages spoken in that region. To address this issue, cross-lingual retrieval attempts to extract useful information simultaneously from multiple languages (Asai et al., 2021; Li et al., 2024a), leading to visible gains in low-resource languages (cf. Figure 1, left).
Figure 1: Illustration of retrieval strategies in mRAG with their performance on the MKQA dataset.
@@
In this paper, we conduct an in-depth assessment of these abilities, using standard accuracy evaluation as well as feature attribution analysis (Figure 2). Unlike recent mRAG evaluations (Chirkova et al., 2024; Park & Lee, 2025), which test the LLM performance for each language in the entire RAG pipeline (i.e., retrieval + generation), we disentangle these two factors and focus on the LLM’s ability to exploit context independently from retrieval quality. Our key contributions include: (1) We evaluate how LLMs leverage retrieved passages in different languages in various multilingual QA tasks, revealing remarkably robust input understanding but much more brittle generation abilities. (2) Besides the standard accuracy evaluation, we apply a recently proposed RAG answer attribution method based on model internals (Qi et al., 2024) to confirm that LLMs consistently incorporate retrieved content from various languages, providing insights from an interpretability perspective. (3) We consider both single-passage and multi-passage mRAG setups and examine how distracting passages in different languages affect model performance, shedding light on the complex interplay between relevance and content of the retrieved passages. Taken together, our results deepen our understanding of how LLMs utilize context in mRAG systems and reveal important areas for future improvements.
Figure 2: To answer our key research question, we propose an evaluation framework composed of two assessment approaches with two different mRAG setups.
@@ 2.1 Retrieval Strategies for mRAG
2.2 Consistency in Multilingual LLMs
Ensuring model consistency across languages is a key objective for multilingual LLMs. A series of recent works has focused on the consistency of factual knowledge encoded in the weights of multilingually pre-trained LLMs (Fierro & Søgaard, 2022; Weber et al., 2023; Qi et al., 2023; Hupkes et al., 2023). Other work has focused on the consistency of domain-specific QA by assessing whether the questions asked by a certain group of people (Schlicht et al., 2025) or about domain-specific knowledge (Yin et al., 2022; Li et al., 2024b) can be correctly answered by LLMs regardless of the query language. Very recently and concurrently with our work, research interest has also risen around the consistency of mRAG pipelines (Wu et al., 2024; Sharma et al., 2024; Park & Lee, 2025).
2.3 Context Utilization in mRAG
Although some studies (Asai et al., 2021; Nie et al., 2023; Chirkova et al., 2024) have demonstrated that cross-lingual retrieval can significantly enhance mRAG answer accuracy, the extent to which LLMs can utilize multilingual contexts consistently remains poorly understood, motivating the present work. The concept of context utilization is also not always clearly defined. Recent and concurrent studies (Wu et al., 2024; Sharma et al., 2024; Park & Lee, 2025) use performance of the complete mRAG pipeline to study context utilization and find that models tend to prefer passages in the query language or Latin scripts. In this paper, we further distinguish between input understanding and decoding capability as key abilities of an mRAG generator, and disentangle them through our experiments, while strictly controlling for retrieval quality.
3 Experimental Setup
3.1 Datasets
XQUAD (Artetxe et al., 2020) is an extension of the extractive English QA dataset SQUAD (Rajpurkar et al., 2016), which contains 1190 questions, each provided with a single relevant passage and a gold answer, all translated into 12 languages. While not being originally designed for RAG evaluation, this dataset is the only one allowing us to assess LLMs’ abilities to use the exact same information provided in different languages, simulating an impossible scenario where retrieval works perfectly in all languages. MKQA (Longpre et al., 2021) is an open domain QA dataset covering 10,000 questions across 24 languages [^1] For the three Chinese variants in MKQA, we select Chinese (Simplified) as the studied language in this paper since it is the most widely used. derived from Natural Questions (Kwiatkowski et al., 2019). Removing the questions without any gold answers provided, we work on a total of 6758 paralleled questions in this paper. Global-MMLU or GMMLU (Singh et al., 2024) is a large multilingual extension of MMLU (Hendrycks et al., 2020) obtained by translating the English instances into 41 languages. Like MMLU, it contains 14042 multi-choice questions that are used to test LLMs’ understanding capability across a range of subjects, like social sciences or medical questions. Each question is provided with four options to choose from. Question examples for all datasets are given in Appendix A.
@@ 3.2 Retrieval and Filtering
Then, we tag the gold passages in $P_{q}$ based on whether they contain the gold answer as a substring, following previous work (Liu et al., 2024a; b). In our experiments, we only consider queries for which $P_{q}$ contains at least one gold passage in any of the studied languages, see resulting # Queries (w/ Gold Pass.) in Table 1. While it may be possible to expand this subset by retrieving more than 30 top passages or by improving retriever quality (Chirkova et al., 2024), we believe our setup is appropriate to study LLMs’ ability to use a variety of multilingual context types that are representative of competitive cross-lingual retrieval results. [^4] Although a large portion of GMMLU queries are filtered out, we argue that the remaining 4136 queries are numerous enough to ensure a robust evaluation. We also verify the diversity of this subset and find a total of 55 covered subjects. See Appendix C for more details on the question subjects and categories.
@@ 3.3 Evaluation Metrics
For XQUAD and MKQA, we follow previous work (Asai et al., 2021; Chirkova et al., 2024) and score answers by strict lexical matching, that is 1 if the entire gold answer string $a^{\ell}$ is a substring of the model response $\mathcal{M}(q^{\ell})$, or 0 otherwise. Since models in mRAG setups often generate the correct answer in the wrong passage language (Chirkova et al., 2024; Zhang et al., 2024), we also measure the proportion of model answers that contain a gold answer in language $\ell^{\prime}$ ($a^{\ell^{\prime}},\ell^{\prime}\neq\ell$).
@@ 3.4 Models
We evaluate four top-performing multilingual LLMs belonging to different model families, which have been used in recent mRAG evaluations (Wang et al., 2024; Thakur et al., 2024), namely: Aya-Expanse-8B (Dang et al., 2024), Llama-3.2-3B-Instruct (Dubey et al., 2024), Gemma-2-9B-it (Team, 2024), and Qwen2.5-7B-Instruct (Yang et al., 2024). Although these models do not officially support some of our studied languages, evidence has shown that LLMs can generalize successfully to unseen languages due to the leak of training data or shared representations (Qi et al., 2023; Budnikov et al., 2024; Lu & Koehn, 2024), which we also observed in preliminary experiments.
@@ 4 Single-Passage mRAG
In XQUAD, where gold passages are translated into 12 languages, we iterate over the 11 out-language passage versions for each query and report the average accuracy. We also report accuracy for the passage language that yielded the best (or worst) answer accuracy overall for each query language. By contrast, the gold passages in MKQA and GMMLU are retrieved from a Wikipedia corpus as explained in Section 3.2, and are not parallel across languages. As a solution, for each query $q^{\ell}$, we randomly sample 3 different out-language passages from $P_{q}$ and report accuracy averaged over the 3 single-passage answers. To maximize the chances of obtaining a model response in the query language $\ell$, we explicitly mention $\ell$ in the instruction, which is itself translated into $\ell$, following Chirkova et al., 2024; Zhang et al., 2024. The detailed prompts are listed in Appendix D.
@@ 4.1 Accuracy Results
Results averaged across all query languages are given in Figure 3, while the full language-specific results are given in Appendix F.
@@
We recall that XQUAD is a distinct dataset, originally developed to evaluate extractive QA, rather than open-domain RAG systems. Nevertheless, it is the only dataset where the exact same gold passage is available in different languages, allowing us to isolate the effect of a passage’s language from that of its content. As shown in Figure 3, providing the gold passage in any language strongly improves answer accuracy compared to the no-context baseline, which is likely due in part to the extractive nature of QA in this dataset. Looking at the passage language, however, we find that in-language passages yield considerably higher accuracy than all out-language settings, including out-language (Best). Moreover, a notable portion of questions are answered correctly but in the wrong language even though the models were explicitly prompted to answer in the target language, which is in line with previous findings (Wu et al., 2024; Chirkova et al., 2024). Even when considering these cases, a visible gap remains between in-language and out-language accuracy across the board on XQUAD. We further analyze this gap through manual error analysis and find that missed matches are often due to the use of synonyms or slight paraphrases of the gold answer, or –in the case of languages with different scripts– to transliteration variations (Knight & Graehl, 1997). See Appendix E for more details.
@@
Figure 4 gives a detailed view of how answer accuracy varies with the language of the provided gold passage. [^5] Here we only consider answers in the correct language, see Appendix F for language-specific accuracies when considering the wrong generation language. As expected, the highest accuracy is always achieved when the retrieved passage is in the same language as the query. Concurrent work (Sharma et al., 2024; Park & Lee, 2025) suggested that models may prefer passage languages that use the same script as the query language, based on a few languages. Because script similarity is a very coarse-grained measure of language similarity that is not informative for many of our language pairs, we turn to finer-grained measures that were previously shown to strongly correlate with cross-lingual consistency of model answers in non-RAG setups (Qi et al., 2023). In particular, we adopt subword vocabulary overlap computed on a reference parallel corpus [^6] Following Qi et al., 2023, we extract the vocabularies from FLORES-200 (Costa-jussà et al., 2022), a strictly parallel corpus covering 200 languages, and measure their pairwise overlap via Jaccard index (Jaccard, 1912). as this was shown to correlate better with response consistency than various typological similarity measures. We compute Pearson and Spearman correlations between subword overlap and answer accuracy for each query language (excluding the case where query and passage are in the same language), however all correlations are low and not statistically significant. Looking back at Figure 4, we notice that shading (i.e., answer accuracy) is relatively consistent within each row, especially on Gemma and Qwen, more so than within each column. In other words, the query language is much more predictive of accuracy than the passage language, suggesting that generating in the target language is the major bottleneck in our setup, which could dominate, if not hide, the effect of similarity with the passage language. Interestingly, we observe that when fed with passages in Thai, which is not officially supported by Aya-Expanse-8B, the model always outperforms the baseline where no context is provided for queries in each language (cf. No-context accuracies in Table 10). This suggests that even though the passages are written in a language that is unseen in the pre-training phase, LLMs may be able to utilize them.

### v3

@@ Abstract
Retrieval-augmented generation (RAG) with large language models (LLMs) has demonstrated strong performance in multilingual question-answering (QA) tasks by leveraging relevant passages retrieved from corpora. In multilingual RAG (mRAG), the retrieved passages can be written in languages other than that of the query entered by the user, making it challenging for LLMs to effectively utilize the provided information. Recent research suggests that retrieving passages from multilingual corpora can improve RAG performance, particularly for low-resource languages. However, the extent to which LLMs can leverage different kinds of multilingual contexts to generate accurate answers, independently from retrieval quality, remains understudied. In this paper, we conduct an extensive assessment of LLMs’ ability to (i) make consistent use of a relevant passage regardless of its language, (ii) respond in the expected language, and (iii) focus on the relevant passage even when multiple ‘distracting’ passages in different languages are provided in the context. Our experiments with four LLMs across three QA datasets covering 48 languages reveal a surprising ability of LLMs to extract relevant information from passages in a different language than the query, but a much weaker ability to produce a full answer in the correct language. Our analysis, based on both accuracy and feature attribution techniques, further shows that distracting passages negatively impact answer quality regardless of their language. However, distractors in the query language exert a slightly stronger influence. Taken together, our findings deepen the understanding of how LLMs utilize context in mRAG systems, providing directions for future improvements.  [^1] All codes and data released at https://github.com/Betswish/mRAG-Context-Consistency.
On the Consistency of Multilingual Context Utilization
in Retrieval-Augmented Generation
Jirui Qi1     Raquel Fernández2     Arianna Bisazza1 1Center for Language and Cognition (CLCG), University of Groningen 2Institute for Logic, Language and Computation (ILLC), University of Amsterdam {j.qi, a.bisazza}@rug.nl, raquel.fernandez@uva.nl
@@ 1 Introduction
Retrieval-augmented generation has shown strong results in multilingual question-answering (QA) tasks (Chirkova et al., 2024; Thakur et al., 2024). Given a query in the user language, informative passages are retrieved from a reference corpus and provided jointly with the query, promoting the large language model (LLM) to generate more precise responses (Lewis et al., 2020; Asai et al., 2021). In multilingual RAG (mRAG), retrieval can be performed either monolingually or cross-lingually. In the former, retrieval is performed only over passages in the same language as the query (Asai et al., 2023; Gao et al., 2023; Fan et al., 2024), which can be successful for high-resource languages. However, this approach is marginally useful, or even harmful, when the question is posed in a low-resource language, since relevant information is likely to be available only in different languages (Muller et al., 2023). In addition, for questions regarding a specific geographical region or culture, essential information may be present only in corpora of the languages spoken in that region. To address this issue, cross-lingual retrieval attempts to extract useful information simultaneously from multiple languages (Asai et al., 2021; Li et al., 2024), leading to visible gains in low-resource languages (Chirkova et al., 2024)
Figure 1: Illustration of the contributions and proposed assessment frameworks of this paper.
@@
In this paper, we conduct an in-depth assessment of these abilities, using standard accuracy evaluation as well as feature attribution analysis. Unlike recent mRAG evaluations (Chirkova et al., 2024; Park and Lee, 2025), which test the LLM performance for each language in the entire RAG pipeline (i.e., retrieval + generation), we disentangle these two factors and focus on the LLM’s ability to exploit context independently from retrieval quality. As shown in Figure 1, our key contributions include: (a) We evaluate how LLMs leverage retrieved passages in different languages in various multilingual QA tasks, revealing remarkably robust input understanding but much more brittle generation abilities. (b) Besides the standard accuracy evaluation, we apply a recently proposed RAG answer attribution method based on model internals (Qi et al., 2024) to confirm that LLMs consistently incorporate retrieved content from various languages, providing insights from an interpretability perspective. (c) We consider both single-passage and multi-passage mRAG setups and examine how distracting passages in different languages affect model performance, shedding light on the complex interplay between relevance and content of the retrieved passages. Taken together, our results deepen our understanding of how LLMs utilize context in mRAG systems and reveal important areas for future improvements.
@@ 2.1 Retrieval Strategies for mRAG
2.2 Consistency in Multilingual LLMs
Ensuring model consistency across languages is a key objective for multilingual LLMs. A series of recent works has focused on the consistency of factual knowledge encoded in the weights of multilingually pre-trained LLMs (Fierro and Søgaard, 2022; Weber et al., 2023; Qi et al., 2023; Hupkes et al., 2023). Other work has focused on the consistency of domain-specific QA by assessing whether the questions asked by a certain group of people (Schlicht et al., 2025) or about domain-specific knowledge (Yin et al., 2022; Li et al., 2025) can be correctly answered by LLMs regardless of the query language. Very recently and concurrently with our work, research interest has also risen around the consistency of mRAG pipelines (Wu et al., 2024; Sharma et al., 2024; Park and Lee, 2025).
2.3 Context Utilization in mRAG
Although some studies (Asai et al., 2021; Nie et al., 2023; Stap and Monz, 2023; Chirkova et al., 2024) have demonstrated that cross-lingual retrieval can significantly enhance mRAG answer accuracy, the extent to which LLMs can utilize multilingual contexts consistently remains poorly understood, motivating the present work. The concept of context utilization is also not always clearly defined. Recent and concurrent studies (Wu et al., 2024; Sharma et al., 2024; Park and Lee, 2025) use performance of the complete mRAG pipeline to study context utilization and find that models tend to prefer passages in the query language or Latin scripts. In this paper, we further distinguish between input understanding and decoding capability as key abilities of an mRAG generator, and disentangle them through our experiments, while strictly controlling for retrieval quality.
3 Experimental Setup
3.1 Datasets
XQUAD (Artetxe et al., 2020) is an extension of the extractive English QA dataset SQUAD (Rajpurkar et al., 2016), which contains 1190 questions, each provided with a single relevant passage and a gold answer, all translated into 12 languages. While not being originally designed for RAG evaluation, this dataset is the only one allowing us to assess LLMs’ abilities to use the exact same information provided in different languages, simulating an impossible scenario where retrieval works perfectly in all languages. MKQA (Longpre et al., 2021) is an open domain QA dataset covering 10,000 questions across 24 languages derived from Natural Questions (Kwiatkowski et al., 2019). Removing the questions without any gold answers provided, we work on a total of 6758 paralleled questions in this paper. Global-MMLU or GMMLU (Singh et al., 2024) is a large multilingual extension of MMLU (Hendrycks et al., 2020) obtained by translating the English instances into 41 languages. Like MMLU, it contains 14042 multi-choice questions that are used to test LLMs’ understanding capability across a range of subjects, like social sciences or medical questions. Each question is provided with four options to choose from. Question examples for all datasets are given in Appendix A.
@@ 3.2 Retrieval and Filtering
Then, we tag the gold passages in $P_{q}$ based on whether they contain the gold answer as a substring, following previous work (Liu et al., 2024, 2025). In our experiments, we only consider queries for which $P_{q}$ contains at least one gold passage in any of the studied languages, see resulting # Queries (w/ Gold Pass.) in Table 1. While it may be possible to expand this subset by retrieving more than 30 top passages or by improving retriever quality (Chirkova et al., 2024), we believe our setup is appropriate to study LLMs’ ability to use a variety of multilingual context types that are representative of competitive cross-lingual retrieval results. [^4] Although a large portion of GMMLU queries are filtered out, we argue that the remaining 4136 queries are numerous enough to ensure a robust evaluation. We also verify the diversity of this subset and find a total of 55 covered subjects. See Appendix C for details on the question subjects and categories.
@@ 3.3 Evaluation Metrics
For XQUAD and MKQA, we follow previous work (Asai et al., 2021) and score answers by strict lexical matching, that is, 1 if the entire gold answer string $a^{\ell}$ is a substring of the model response $\mathcal{M}(q^{\ell})$, or 0 otherwise. Since models in mRAG setups often generate the correct answer in the wrong passage language (Chirkova et al., 2024; Zhang et al., 2024), we also measure the proportion of model answers that contain a gold answer in language $\ell^{\prime}$ ($a^{\ell^{\prime}},\ell^{\prime}\neq\ell$).  [^5] Since we focus on the language of model responses and outright cross-language generation (i.e., whether the gold answer appears in a different language) where small orthographic variants can be decisive, particularly for phonologically similar languages, we do not adopt the variant of the softer lexical metric (Chirkova et al., 2024) (3-gram recall), which tolerates minor orthographic differences and could blur the distinctions. Nevertheless, as exact matching could be overly strict, we further adopt two complementary metrics (BERTScore and GPT-4.1-nano) on XQUAD. Similar results are observed, providing more insights and enhancing the robustness of our analysis. See Appendix D for more details.
@@ 3.4 Models
We evaluate four top-performing multilingual LLMs belonging to different model families, which have been used in recent mRAG evaluations (Wang et al., 2024; Thakur et al., 2024), namely: Aya-Expanse-8B (Dang et al., 2024), Llama-3.2-3B-Instruct (Dubey et al., 2024), Gemma-2-9B-it (Team, 2024), and Qwen2.5-7B-Instruct (Yang et al., 2024). Although these models do not officially support some of our studied languages, evidence has shown that LLMs can generalize successfully to unseen languages due to the leak of training data or shared representations (Qi et al., 2023; Budnikov et al., 2024; Lu and Koehn, 2024), which we also observed in preliminary experiments.
@@ 4 Single-Passage mRAG
In XQUAD, where gold passages are translated into 12 languages, we iterate over the 11 out-language passage versions for each query and report the average accuracy. We also report accuracy for the passage language that yielded the best (or worst) answer accuracy overall for each query language. By contrast, the gold passages in MKQA and GMMLU are retrieved from a Wikipedia corpus as explained in Section 3.2, and are not parallel across languages. As a solution, for each query $q^{\ell}$, we randomly sample 3 different out-language passages from $P_{q}$ and report accuracy averaged over the 3 single-passage answers. To maximize the chances of obtaining a model response in the query language $\ell$, we explicitly mention $\ell$ in the instruction, which is itself translated into $\ell$, following Chirkova et al., 2024; Zhang et al., 2024. The detailed prompts are listed in Appendix E.
@@ 4.1 Accuracy Results
Results averaged across all query languages are given in Figure 2, while the full language-specific results are given in Appendix I.
@@
We recall that XQUAD is a distinct dataset, originally developed to evaluate extractive QA, rather than open-domain RAG systems. Nevertheless, it is the only dataset where the exact same gold passage is available in different languages, allowing us to isolate the effect of a passage’s language from that of its content. As shown in Figure 2, providing the gold passage in any language strongly improves answer accuracy compared to the no-context baseline, which is likely due in part to the extractive nature of QA in this dataset. Looking at the passage language, however, we find that in-language passages yield considerably higher accuracy than all out-language settings, including out-language (Best). Moreover, a notable portion of questions are answered correctly but in the wrong language even though the models were explicitly prompted to answer in the target language, which is in line with previous findings (Wu et al., 2024; Chirkova et al., 2024). Even when considering these cases, a visible gap remains between in-language and out-language accuracy across the board on XQUAD. We further analyze this gap through manual error analysis and find that missed matches are often due to the use of synonyms or slight paraphrases of the gold answer, or –in the case of languages with different scripts– to transliteration variations (Knight and Graehl, 1997). See Appendix F for more details.
@@
Figure 3 gives a detailed view of how answer accuracy varies with the language of the provided gold passage. [^6] Here we only consider answers in the correct language, see Appendix I for language-specific accuracies when considering the wrong generation language. As expected, the highest accuracy is always achieved when the retrieved passage is in the same language as the query. Concurrent work (Sharma et al., 2024; Park and Lee, 2025) suggested that models may prefer passage languages that use the same script as the query language, based on a few languages. Because script similarity is a very coarse-grained measure of language similarity that is not informative for many of our language pairs, we turn to finer-grained measures that were previously shown to strongly correlate with cross-lingual consistency of model answers in non-RAG setups (Qi et al., 2023). In particular, we adopt subword vocabulary overlap computed on a reference parallel corpus [^7] Following Qi et al., 2023, we extract the vocabularies from FLORES-200 (Costa-jussà et al., 2022), a strictly parallel corpus covering 200 languages, and measure their pairwise overlap via Jaccard index (Jaccard, 1912). as this was shown to correlate better with response consistency than various typological similarity measures. We compute Pearson and Spearman correlations between subword overlap and answer accuracy for each query language (excluding the case where query and passage are in the same language), however all correlations are low and not statistically significant. Looking back at Figure 3, we notice that shading (i.e., answer accuracy) is relatively consistent within each row, especially on Gemma and Qwen, more so than within each column. In other words, the query language is much more predictive of accuracy than the passage language, suggesting that generating in the target language is the major bottleneck in our setup, which could dominate, if not hide, the effect of similarity with the passage language.
Additionally, we also investigate if advanced prompts with multi-step instructions (refer to Appendix G) or larger models (open-source Gemma3-27B-IT and closed-source GPT-5-nano estimated at 8-18B parameters [^8] https://www.r-bloggers.com/2025/08/how-many-parameters-does-gpt-5-have/; Appendix H) can mitigate the language mismatch issue in model answers. However, the problem persists, further reinforcing our finding that multilingual RAG systems face an inherent decoding limitation. Interestingly, we observe that when fed with passages in Thai, which is not officially supported by Aya-Expanse-8B, the model always outperforms the baseline where no context is provided for queries in each language (cf. No-context accuracies in Table 13). This suggests that even though the passages are written in a language that is unseen in the pre-training phase, LLMs may be able to utilize them.